• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 08:03:27 on February10,2002

    Entry number 82483

    keyword=shift summary

    00:00, controlled access, Bob is working on T4;
    00:35, hall in beam permit;
    00:43, turn on deflection magnet, current 600A;

    Pending: from main control screen, LeCroy HV gui has communication problem.

    00:57, Run 2376, 15cm Empty target with radiator, RCS magnet on, kinematics 5Bf*, stopped due to wrong prescale factors;
    01:05~01:33, Run 2377, kinematics 5B(f*), good run;

    01:35, move target to LH2+radiator, start kinematics 5B(e);
    01:40~01:42, Run 2378, 15uA, check prescale factor;
    01:44, start production run with #2379, beam current 25uA;
    03:31, reduce target loop3 coolant flow (JT valve) to 40%;
    04:40, ROC4 crashed, run#2384 end of run failed. Reboot ROC4, kcoda, restart runcontrol.
    Run#2385 is junk, #2386 end failed;
    05:10, ROC4 crashed again, reboot ROC4, kcoda, reduce beam current to 20 uA, reset crate for ROC4;
    05:26, MCC called, problem with cavity(?), will take 10 minutes to fix it.
    05:38, beam back;
    05:46, veto HV tripped;
    07:07, MCC called, beam will be off for 15 minutes;
    07:49, beam back, run#2392 is junk (no beam);
    07:56, run #2393, beam position not stable in the first 3 minutes;
    08:00, run #2393 continue to the next shift.

    ----------------------------
    kine#5B(f*) production runs:
    run#....start~end......#events
    2377....01:05~01:33......100K
    ----------------------------
    kine#5B(e) production runs:
    run#....start~end.........#events... comments
    2379....01:44~02:10.......300K.......good
    2380....02:11~02:34.......300K.......good
    2381....02:35~03:03.......300K.......good
    2382....03:03~03:38.......300K.......good
    2383....03:38~04:00.......300K.......good
    2384....04:01~04:40..........?...........end failed
    2386....04:42~04:54........69K........end failed
    2387....04:54~05:10.......154K.......end failed
    2388....05:13~05:26.........81K.......good
    2389....05:42~06:16.......300K.......good
    2390....06:17~07:00.......300K.......good
    2391....07:00~07:09.........17K.......good
    2393....07:56~......
    -----------------------------

    Pending: from main control screen, LeCroy HV gui has communication problem.